# Proper maps of annuli

### v1

@@ Abstract.
We study rational proper holomorphic maps of annuli in complex euclidean spaces, that is, domains with $U(n)$ as the automorphism group. By the Hartogs phenomenon and a result of Forstnerič, such maps are always rational and extend to proper maps of balls. We first prove that a proper map of annuli from $n$ dimensions to $N$ dimensions where $N<\binom{n+1}{2}$ is always an affine embedding. This inequality is sharp as the homogeneous map of degree 2 satisfies $N=\binom{n+1}{2}$. Next we find a necessary and sufficient condition for a map to be homogeneous: A proper map of annuli is homogeneous if and only if its general hyperplane rank, the affine dimension of the image of a general hyperplane, is exactly N1N-1. This result is really a result classifying homogeneous proper maps of balls. A homogeneous proper ball map takes spheres centered at the origin to spheres centered at the origin. We show that if a proper ball map has general hyperplane rank N1N-1 and takes one sphere centered at the origin to a sphere centered at the origin, then it is homogeneous. Another corollary of this result is a complete classification of proper maps of annuli from dimension 2 to dimension 3.
@@ 1. Introduction
Let $\mathbb{B}_{n}\subset\mathbb{C}^{n}$ denote the unit ball. A well-studied problem in complex analysis is to attempt to classify the proper holomorphic maps $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$. If $n=N=1$, then by a classical theorem of Fatou [fatou-1923-fonctions], $f$ is a finite Blaschke product and hence rational. If $n=N>1$, then by a theorem of Alexander [alexander-1977-proper], $f$ is an automorphism of the unit ball. No maps exist if $N<n$ for elementary reasons. Dor and Løw [dor-1990-proper, low-1985-embeddings] show that if $N>n>1$, there exist non-rational maps, even in the case of codimension 1, that is, $N=n+1$. If in the case $n>1$ one assumes some regularity on the boundary, then by a theorem of Forstnerič [forstneric-1989-extending], all such $f$ are rational with the degree bounded by an integer depending on $n$ and $N$. Faran [faran-1982-maps] showed that proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$ that are $C^{3}$ up to the boundary are equivalent to one of four polynomial maps. For maps that are $C^{3}$ up to the boundary, Webster [webster-1979-mapping] showed that for $n\geq 3$, there is no map from $\mathbb{B}_{n}$ to $\mathbb{B}_{n+1}$ that is not equivalent up to automorphisms (spherically equivalent) to the linear embedding $z\oplus 0$. Faran [faran-1986-linearity] proved the same conclusion (only the linear embedding) for maps real-analytic up to the boundary for all $N$ up to N=2n2N=2n-2. Huang [huang-1999-linearity] improved the regularity to $C^{2}$. For N=2n1N=2n-1, the Whitney map $(z^{\prime},z_{n})\mapsto(z^{\prime},z_{n}\otimes z)$ is not spherically equivalent to the linear embedding. The map $z^{\otimes d}$ is a proper map of $\mathbb{B}_{n}$ to $\mathbb{B}_{N}$. This map can be symmetrized (as $z^{\otimes d}$ contains some duplicate entries) via a unitary mapping to what we will call $H_{d}$ direct sum a number of zero components. Rudin [rudin-1984-homogeneous] proved that every homogeneous proper ball map is, up to a unitary, $H_{d}\oplus 0$. See D’Angelo [dangelo-1988-polynomial] for a simpler proof of this fact. In general, see the books by D’Angelo [dangelo-1993-several, dangelo-2019-hermitian, dangelo-2021-rational] for a more thorough treatment of proper maps of balls.
@@
The ball can be regarded as the strongly pseudoconvex domain with the most symmetry. Let us consider then a non-pseudoconvex domain with a lot of symmetry, the annulus. Write $\mathbb{A}_{n,r}=\mathbb{B}_{n}\setminus\overline{r\mathbb{B}_{n}}=\{z\in\mathbb{C}^{n}:r<\mathopen{}\mathclose{{\left\lVert z}}\right\rVert<1\}$ for an annulus, where we will always assume that $0<r<1$. The automorphism group of $\mathbb{A}_{n,r}$ is the unitary group as long as $n>1$. When $n=1$ there are also the automorphisms $e^{i\theta}\frac{r}{z}$. The same classical argument to classify the automorphism group $\mathbb{A}_{1,r}$ can be extended to show that in one dimension a proper holomorphic map $f\colon\mathbb{A}_{1,r}\to\mathbb{A}_{1,R}$ exists only if $R=r^{d}$ for some $d\in\mathbb{N}$ in which case it is either equal to $e^{i\theta}z^{d}$ or $r^{d}e^{i\theta}z^{-d}$. In dimension $n\geq 2$, due to the inner pseudoconcave boundary, via the Hartogs phenomenon and the theorem of Forstnerič on rationality of proper ball maps, all proper holomorphic maps $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ are rational, see Theorem˜2.2, but they can be more complicated. In fact, holomorphic proper maps of annuli are in one to one correspondence with rational proper maps of balls that take a sphere centered at zero to another sphere centered at zero. See [helal-2025-proper] for more details.
@@
For $z\in\mathbb{C}^{n}$, if $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$ is constant, then
(1) $\mathopen{}\mathclose{{\left\lVert H_{d}(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z^{\otimes d}}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2d}$
@@
is constant. Hence, $H_{d}$ is a map that takes any sphere of radius $r$ centered at the origin to a sphere of radius $r^{d}$ centered at the origin. Moreover, it is not difficult to show that it takes $\mathbb{A}_{n,r}$ (for any $0<r<1$) properly to $\mathbb{A}_{N,R}$, where $N$ is the number of components of $H_{d}$ and $R=r^{d}$. Similarly, if we have any two proper maps $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{1},R_{1}}$ and $G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{2},R_{2}}$ and $t\in(0,1)$, then $\sqrt{1-t}\,F\oplus\sqrt{t}\,G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper map where $N=N_{1}+N_{2}$ and $R^{2}=(1-t)R_{1}^{2}+tR_{2}^{2}$. We call such a weighted direct sum a juxtaposition. So juxtapositions of proper maps of annuli are proper maps of annuli as long as the source annulus is the same. Based on the one-dimensional result, noting that $z^{\otimes d}$ is the correct multidimensional analogue of $z^{d}$, one could (incorrectly) conjecture that all such proper maps are juxtapositions of homogeneous maps. However, other maps do exist, see Example˜2 and more generally [helal-2025-proper].
@@
Our goal is to start a classification of proper maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$. We say $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ and $g\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ are unitarily equivalent (or simply equivalent) if there exist unitary matrices $V\in U(n)$ and $W\in U(N)$ so that
@@
As we noted, $U(n)$ is the holomorphic automorphism group of $\mathbb{A}_{n,r}$. If $n=N>1$, then a proper map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ exists only if $r=R$ and $f$ is unitarily equivalent to the identity $z$, see Corollary˜2.6. If $N>n$, then an affine embedding exists whenever $r\leq R<1$, but is not necessarily $z\oplus 0$ due to the different automorphism group. In particular, $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree 1, then $f$ is unitarily equivalent to the affine embedding
@@
where the last 0 has Nn1N-n-1 terms. That is, we shift the linear embedding by adding the proper constant in the $(n+1)$th component. See Proposition˜2.7.
As we noted above, for sufficiently smooth proper maps between balls, $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ no maps other than the linear embedding exist for N<2n1N<2n-1, and then when N=2n1N=2n-1, the Whitney mapping exists and is not spherically equivalent to a linear embedding. Curiously, for annulus maps, this gap is much longer and goes until $N=\binom{n+1}{2}$.
@@
Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map and $2\leq n<N<\binom{n+1}{2}$. Then $f$ is unitarily equivalent to an affine embedding (3) and $r\leq R<1$. The bound on $N$ is sharp as the homogeneous map $H_{2}$ takes the annulus $\mathbb{A}_{n,r}$ properly to $\mathbb{A}_{\binom{n+1}{2},r^{2}}$.
As homogeneous maps play an outsize role in the study of proper maps of annuli we next find a property that determines a map to be homogeneous. The framework that we find applies more broadly to proper maps of balls and so we state it first in that setting. First we define a certain invariant for proper mappings that has, in various guises, appeared in the study of ball maps. First let $f\colon U\subset\mathbb{C}^{n}\to\mathbb{C}^{N}$ be any holomorphic map. If $X\subset\mathbb{C}^{N}$, write $\operatorname{aff}(X)$ for the affine hull of $X$, that is, the smallest affine subspace of $\mathbb{C}^{N}$ containing $X$. Define
@@
We call $k_{f}$ the general hyperplane rank. One could think of $k_{f}$ as a certain measure of complexity. It is not difficult to see that $k_{f}$ is attained on a generic hyperplane, and it is in general quite easy to compute, especially for rational maps. While for an arbitrary holomorphic map one generally expects $k_{f}=N$, via a Segre-variety argument, a rational proper map of balls has $k_{f}\leq N-1$. The idea that Dor [dor-1990-proper] used to show a proper map is not rational is to, in other words, show that the constructed map has $k_{f}=N$ as any rational ball map has $k_{f}=N-1$. If $k_{f}=N-1$, then we obtain a map of Grassmannians. That technique was used for example by Faran to prove his classification of maps from $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$. The number $k_{f}$ is related to the $X$-variety that Forstnerič used to show the rationality result. In particular, for a rational map, if the dimension of the generic fiber of the $X$-variety is $\ell$, then $k_{f}=N-\ell-1$. D’Angelo [dangelo-2003-Xvariety] has studied the $X$-variety for rational maps and developed a technique of quickly computing fibers of the $X$-variety as the nullity of a certain matrix $C(w)$. The nullity of $C(w)$ for a generic $w$ is then the $\ell$. A related concept is the embedding dimension $N_{f}$, that is, the dimension of the smallest affine subspace $S$ containing $f(\mathbb{B}_{n})$ (or $f(\mathbb{A}_{n,r})$). In other words, $N_{f}=\dim\operatorname{aff}f(\mathbb{B}_{n})$. In either case, restricting to $S$ on the target we obtain a proper map between balls or annuli with target dimension $N_{f}$. While for rational proper maps, often $k_{f}=N_{f}-1$, it is not always true. Example˜2 mentioned above has $k_{f}=N_{f}-3$. A juxtaposition of two rational maps also has $k_{f}\leq N_{f}-2$, see Proposition˜2.12. What we prove is that for proper maps of annuli, $k_{f}=N_{f}-1$ is precisely the condition that characterizes homogeneous maps $H_{d}$.
@@
Suppose $n\geq 2$ and $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a holomorphic proper map. Then $k_{f}=N-1$ if and only if $N=\binom{n+d-1}{d}$, $R=r^{d}$, and
@@
One direction of this theorem is obvious, it is that $k_{f}=N-1$ implies that $f$ is homogeneous that is remarkable. The proof relies on a result from [helal-2025-proper], where the authors showed that a rational map that takes infinitely many spheres centered at the origin to spheres centered at the origin is a juxtaposition of homogeneous maps. We note that if we wish to state the theorem in terms of $N_{f}$, then we ought to say that $k_{f}=N_{f}-1$ if and only if $f=U(H_{d}\oplus 0)$. We note that $k_{f}=N-1$ implies that $N=N_{f}$ for any proper rational map of annuli (or balls).
@@
Suppose $n\geq 2$ and $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ is a rational proper map such that $k_{f}=N-1$ and $f(rS^{2n-1})\subset RS^{2N-1}$ for some $r,R\in(0,1)$. Then there exists a unitary map $U$ and an integer $d$ such that $f=UH_{d}$.
The proof follows by using the $k_{f}=N-1$ condition to show that $f$ must in fact take infinitely many spheres centered at zero to spheres centered at zero.
@@
Let us give the organization of this paper. In Section˜2, we state and prove the basic results for maps of annuli. In Section˜3, we prove the first gap theorem, that is, Theorem˜1.1. In Section˜4, we prove the classification of homogeneous maps in terms of $k_{f}$, that is, Theorem˜1.2. We also prove the classification of annuli maps in dimension 2 and 3, that is, Theorem˜1.4. Finally, in Section˜5, we give some further remarks and open questions.

### v2

@@ Abstract.
We study proper holomorphic maps of annuli in complex Euclidean spaces, that is, domains with $U(n)$ as the automorphism group. By the Hartogs phenomenon and a result of Forstnerič, such maps are always rational and extend to proper maps of balls. We first prove that a proper map of annuli from $n$ dimensions to $N$ dimensions where $N<\binom{n+1}{2}$ is always an affine embedding. This inequality is sharp as the homogeneous map of degree 2 satisfies $N=\binom{n+1}{2}$. Next we find a necessary and sufficient condition for a map to be homogeneous: A proper map of annuli is homogeneous if and only if its general hyperplane rank, the affine dimension of the image of a general hyperplane, is exactly N1N-1. As a corollary, we obtain a classification of homogeneous proper maps of balls. A homogeneous proper ball map takes all spheres centered at the origin to spheres centered at the origin. We show that if a proper ball map has general hyperplane rank N1N-1 and takes one sphere centered at the origin to a sphere centered at the origin, then it is homogeneous. Another corollary of this result is a complete classification of proper maps of annuli from dimension 2 to dimension 3. Finally, we give a complete normal form of rational proper maps of annuli of degree 2.
@@ 1. Introduction
One of the fundamental questions in complex analysis is to understand proper holomorphic maps $f\colon U\to V$ between two domains $U\subset\mathbb{C}^{n}$ and $V\subset\mathbb{C}^{N}$. In general, few if any such maps exist unless $U$ and $V$ are related. If $U$ and $V$ are the same type of domains and have many symmetries, then many such maps exist in general. The domain with the “largest” automorphism group is the unit ball $\mathbb{B}_{n}\subset\mathbb{C}^{n}$ and the space of proper maps of balls $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ has a rich structure and has been extensively studied. We will focus on proper maps of annuli, as annuli are domains with a large automorphism group. Write $\mathbb{A}_{n,r}=\mathbb{B}_{n}\setminus\overline{r\mathbb{B}_{n}}=\{z\in\mathbb{C}^{n}:r<\mathopen{}\mathclose{{\left\lVert z}}\right\rVert<1\}$ for an annulus, where we will always assume that $0<r<1$. The automorphism group of $\mathbb{A}_{n,r}$ is the unitary group as long as $n>1$. When $n=1$ there are also the automorphisms $e^{i\theta}\frac{r}{z}$. Notice that unlike the ball, the annulus is non-pseudoconvex; its boundary consists of a strongly pseudoconvex part and a strongly pseudoconcave part.
It is not difficult to show that if $n>1$, then proper holomorphic maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ extend to proper maps of balls that also take the $r$-sphere to the $R$-sphere. Their study therefore involves proper maps of balls. Our goal is to prove results for annuli analogous to the known results for balls and to contrast where these results fundamentally differ. We therefore very briefly discuss some of the known results for balls.
@@
Consider then a proper holomorphic map $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$. If $n=N=1$, then by a classical theorem of Fatou [16], $f$ is a finite Blaschke product and hence rational. If $n=N>1$, then by a theorem of Alexander [1], $f$ is an automorphism of the unit ball. Dor [13] and Løw [22] show that if $N>n>1$, there exist nonrational maps, even in the case of codimension 1, that is, $N=n+1$. If $n>1$ and the map is sufficiently smooth up to the boundary, then by a theorem of Forstnerič [17], all such $f$ are rational with the degree bounded by an integer depending on $n$ and $N$. Faran [14] showed that proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$ that are $C^{3}$ up to the boundary are spherically equivalent (up to composition with automorphisms of the balls) to one of four polynomial maps. For $n\geq 3$, by a series of papers of Webster [24], Faran [15], and Huang [20], every proper map from $\mathbb{B}_{n}$ to $\mathbb{B}_{N}$ for $N\leq 2n-2$ that is sufficiently smooth up to the boundary is spherically equivalent to the linear embedding. For N=2n1N=2n-1, the Whitney map $(z^{\prime},z_{n})\mapsto(z^{\prime},z_{n}\otimes z)$ is not spherically equivalent to the linear embedding. Interestingly, this map has no analogue for annuli. In general, see the books by D’Angelo [9, 11, 12] for a more thorough treatment of proper maps of balls.
@@
Let us move to maps of annuli. When $n=1$, the same classical argument to compute the automorphism group of $\mathbb{A}_{1,r}$ extends to show that a proper holomorphic map $f\colon\mathbb{A}_{1,r}\to\mathbb{A}_{1,R}$ exists only if $R=r^{d}$ for some $d\in\mathbb{N}$ in which case $f(z)=e^{i\theta}z^{d}$ or $f(z)=r^{d}e^{i\theta}z^{-d}$. As we said before, if $n>1$, annulus maps extend to ball maps. By Forstnerič, the maps must be rational. See Theorem˜2.2 (proved in [19]) for more details. So when $n>1$, holomorphic proper maps of annuli are in one-to-one correspondence with rational proper maps of balls that take a sphere centered at zero to another sphere centered at zero.
@@
A natural map that is a proper map for both balls and annuli is the homogeneous map $z^{\otimes d}$ (the $d$th tensor power). That is because if $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$ is constant, then
(1) $\mathopen{}\mathclose{{\left\lVert z^{\otimes d}}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2d}$
is constant. This map can be symmetrized (as $z^{\otimes d}$ contains some duplicate entries) via a unitary map to what we will call $H_{d}$ direct sum a number of zero components. In particular, $\mathopen{}\mathclose{{\left\lVert H_{d}(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z^{\otimes d}}}\right\rVert^{2}$. Rudin [23] proved that every homogeneous proper ball map is, up to a unitary, $H_{d}\oplus 0$. See D’Angelo [8, 10, 9] for a much simpler proof of this fact and many more details and results about the homogeneous map.
@@
Hence, $H_{d}$ is a map that takes any sphere of radius $r$ centered at the origin to a sphere of radius $r^{d}$ centered at the origin, and hence takes $\mathbb{A}_{n,r}$ (for any $0<r<1$) properly to $\mathbb{A}_{N,R}$, where $N=\binom{n+d-1}{d}$ is the number of components of $H_{d}$ and $R=r^{d}$. We can now construct new maps by what we will call juxtaposition. Given two proper annulus maps with the same source, $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{1},R_{1}}$ and $G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{2},R_{2}}$ and $t\in(0,1)$, then $\sqrt{1-t}\,F\oplus\sqrt{t}\,G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper map where $N=N_{1}+N_{2}$ and $R^{2}=(1-t)R_{1}^{2}+tR_{2}^{2}$. Based on the one-dimensional result, noting that $z^{\otimes d}$ is the correct multidimensional analogue of $z^{d}$, one could (incorrectly) conjecture that all such proper maps are juxtapositions of homogeneous maps. However, other maps do exist, see Section˜6 and more generally [19].
@@
Our goal is to classify proper holomorphic maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$. We say $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ and $g\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ are unitarily equivalent (or simply equivalent) if there exist unitary matrices $V\in U(n)$ and $W\in U(N)$ so that
@@
As we noted, $U(n)$ is the holomorphic automorphism group of $\mathbb{A}_{n,r}$. If $n=N>1$, then a proper map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ exists only if $r=R$ and $f$ is unitarily equivalent to the identity $z$, see Corollary˜2.7. If $N>n$, then an affine embedding exists whenever $r\leq R<1$, but is not necessarily $z\oplus 0$ due to the different automorphism group. In particular, if $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree 1, then $f$ is unitarily equivalent to the affine embedding
@@
where the last 0 has Nn1N-n-1 terms. That is, we shift the linear embedding by adding the proper constant in the $(n+1)$th component. Allowing for constants in juxtaposition, this map is a juxtaposition of the identity $z$ and the constant map $1$. See Proposition˜2.8.
As we noted above, for sufficiently smooth proper maps between balls, $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ no maps other than the linear embedding exist for N<2n1N<2n-1, and then when N=2n1N=2n-1, the Whitney map exists and is not spherically equivalent to a linear embedding. Curiously, for annulus maps, this gap is much longer and goes until $N=\binom{n+1}{2}$.
@@
Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map and $2\leq n<N<\binom{n+1}{2}$. Then $r\leq R<1$ and $f$ is unitarily equivalent to an affine embedding (3). The bound on $N$ is sharp as the homogeneous map $H_{2}$ takes the annulus $\mathbb{A}_{n,r}$ properly to $\mathbb{A}_{\binom{n+1}{2},r^{2}}$.
As homogeneous maps play a significant role in the study of proper maps of annuli, we next find a property that determines a map to be homogeneous. We define an invariant $k_{f}$ that has appeared in many guises in the study of ball maps. Let $f\colon U\subset\mathbb{C}^{n}\to\mathbb{C}^{N}$ be any map. If $X\subset\mathbb{C}^{N}$, write $\operatorname{aff}(X)$ for the affine hull of $X$, that is, the smallest affine subspace of $\mathbb{C}^{N}$ containing $X$. Define
@@
We call $k_{f}$ the general hyperplane rank. One could think of $k_{f}$ as a certain measure of complexity. For a holomorphic map $f$, $k_{f}$ is attained on a generic hyperplane, and it is in general quite easy to compute, especially for rational maps. While for an arbitrary holomorphic map one generally expects $k_{f}=N$, via a Segre-variety argument, a rational proper map of balls (or annuli) has $k_{f}\leq N-1$. This fact has been used many times in the study of ball maps. The integer $k_{f}$ is also related to the $X$-variety. See Section˜3 for more on this connection and the history. A related concept is the embedding dimension $N_{f}$, that is, the dimension of the smallest affine subspace $S$ containing $f(\mathbb{A}_{n,r})$ (or $f(\mathbb{B}_{n})$). In other words, $N_{f}=\dim\operatorname{aff}f(\mathbb{A}_{n,r})$. In either case, restricting the codomain to $S$ obtains a proper map between annuli (or balls) with target dimension $N_{f}$. Note that it is not always true that $k_{f}=N_{f}-1$. If $f$ is a juxtaposition of two rational maps, then $k_{f}\leq N_{f}-2$, see Proposition˜3.5. What we prove is that for proper maps of annuli, $k_{f}=N_{f}-1$ is precisely the condition that characterizes homogeneous maps $H_{d}$.
@@
Suppose $n\geq 2$ and $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map. Then $k_{f}=N-1$ if and only if $N=\binom{n+d-1}{d}$, $R=r^{d}$, and
@@
One direction of this theorem is obvious. The remarkable direction is that $k_{f}=N-1$ implies that $f$ is homogeneous. The proof relies on a result from [19], where the authors showed that a rational map that takes infinitely many spheres centered at the origin to spheres centered at the origin is a juxtaposition of homogeneous maps. To state the theorem in terms of $N_{f}$, we ought to say that $k_{f}=N_{f}-1$ if and only if $f=U(H_{d}\oplus 0)$. Note that $k_{f}=N-1$ implies $N=N_{f}$ for any proper rational map of annuli (or balls).
@@
Suppose $n\geq 2$ and $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ is a proper holomorphic map such that $k_{f}=N-1$ and $f(rS^{2n-1})\subset RS^{2N-1}$ for some $r,R\in(0,1)$. Then there exists a unitary map $U$ and an integer $d$ such that $f=UH_{d}$.
To prove the theorem, we note that we have a rational map taking two distinct spheres to two distinct spheres. The condition $k_{f}=N-1$ is then used to show that $f$ must in fact take infinitely many spheres to spheres.
@@
It is also interesting to classify maps by degree. Degree one maps are unitaries possibly composed with an affine embedding if the target dimension is larger than the embedding dimension. We completely classify degree two maps in Section˜6. There exist other degree two maps than the homogeneous map $H_{2}$ or a juxtaposition of $H_{2}$ and $H_{1}$ and/or $H_{0}$. For proper ball maps, all degree two maps are spherically equivalent to monomial maps by the work of the second author [21]. For annuli maps where the automorphism group is the unitary, the conclusion is more elaborate. For every $r$ and $R$ with $r^{2}<R<1$, there exists a one-parameter family of rational proper maps from $\mathbb{A}_{n,r}$ to $\mathbb{A}_{N,R}$, where $N_{f}=N=\binom{n+1}{2}+n$ and these maps are not equivalent to a polynomial map.
Let us give the organization of this paper. In Sections˜2 and 3, we state and prove the basic results for maps of annuli and the general hyperplane rank $k_{f}$. In Section˜4, we prove the first gap theorem, that is, Theorem˜1.1. In Section˜5, we prove the classification of homogeneous maps in terms of $k_{f}$, that is, Theorem˜1.2. We also prove the classification of annuli maps in dimension 2 and 3, that is, Theorem˜1.4. In Section˜6, we give a complete classification of second degree rational annuli maps. Finally, in Section˜7, we give some further remarks and open questions.
The authors would like to acknowledge John D’Angelo for many insightful comments on this work, and the second author in particular would like to acknowledge D’Angelo for introducing him to the problems surrounding proper maps of balls and for the many years of extremely fruitful discussions on related topics. The authors would also like to acknowledge Dusty Grundmeier for useful suggestions on this work.